# Stability of Poiseuille Flow of Navier-Stokes Equations on $\mathbb{R}^{2}$

### v1

@@ 3.1 Technical Lemmas
$\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}|k|\|\nabla_{k}\psi_{k}\|_{\infty}^{2}\mathrm{d}k\lesssim\nu^{-\frac{1}{3}}\mathcal{D}.$
@@
$\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}|k|\|\nabla_{k}\psi_{k}\|_{\infty}^{2}\mathrm{d}k\lesssim\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}|k|\|\nabla_{k}\psi_{k}\|_{2}\|\partial_{y}\nabla_{k}\psi_{k}\|_{2}\mathrm{d}k\lesssim\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|\omega_{k}\|_{2}^{2}\mathrm{d}k.$
@@ 3.2 Bound on $T_{1}$
Since $M_{k}(t)$ is uniformly bounded above and below, we will absorb it into the implicit constant.
@@
$|T_{1,LH}|\lesssim\iint\langle c\lambda_{k}t\rangle^{J}\langle k\rangle^{m}\|\omega_{k}\|_{2}\langle c\lambda_{k^{\prime}}t\rangle^{J}\langle k^{\prime}\rangle^{m}\|\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}.$
Then we use Young’s convolution inequality to place the $k$ and $k^{\prime}$ factors into $L_{k}^{2}$ and the $k-k^{\prime}$ factor into $L_{k}^{1}$,
$|T_{1,LH}|$ $\lesssim\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k$
$\lesssim\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}(\gamma_{k}\nu)^{-1}D_{k}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\nu^{-1}D_{k}\mathrm{d}k\right)^{\frac{1}{2}}\int\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k.$
@@
$|T_{1,LH}|\lesssim(\nu^{-\frac{1}{3}}\mathcal{D})^{\frac{1}{2}}(\nu^{-1}\mathcal{D})^{\frac{1}{2}}\mathcal{E}^{\frac{1}{2}}=\nu^{-\frac{2}{3}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
Turning to $T_{1,HL}$, we similarly use Cauchy-Schwarz inequality in $y$, together with Young’s convolution inequality and Lemma 3.4,
$|T_{1,HL}|$ $\lesssim\iint\langle c\lambda_{k}t\rangle^{J}\langle k\rangle^{m}\|\omega_{k}\|_{2}\langle c\lambda_{k-k^{\prime}}t\rangle^{J}\langle k-k^{\prime}\rangle^{m}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{2}\|\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}$
$\lesssim\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|\nabla_{k}\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int\|\nabla_{k}\omega_{k}\|_{\infty}\mathrm{d}k$
$\lesssim\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}(\gamma_{k}\nu)^{-1}D_{k}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}^{-1}E_{k}\mathrm{d}k\right)^{\frac{1}{2}}\nu^{-\frac{5}{6}}\mathcal{D}^{\frac{1}{2}}$
$\lesssim\nu^{-\frac{2}{3}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
3.3 Bound on $T_{2}$
Using integration by parts,
$T_{2}=-c_{\alpha}\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\alpha_{k}\mathrm{Re}\langle\Delta_{k}\omega_{k},NL_{k}\rangle\mathrm{d}k.$
Using the same frequency decomposition as $T_{1}$, we note that $\alpha_{k}\leq\nu^{\frac{2}{3}}$, and the estimates are similar to $T_{1}$. The bounds of $T_{2}$ is $|T_{2}|\lesssim\nu^{-\frac{2}{3}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}$.
3.4 Bound on $T_{3}$
We still use the frequency decomposition $1=1_{|k-k^{\prime}|\leq\frac{|k|}{2}}+1_{|k-k^{\prime}|>\frac{|k|}{2}}$. Recalling the definition of $\beta_{k}$, we obtain $\beta_{k}|k|\leq 1$, and then, similarly to $T_{1}$, we have $|T_{3}|\lesssim\nu^{-\frac{2}{3}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}$.
3.5 Bound on $T_{4}$
We use integration by parts and separate $T_{4}$ into two parts,
$T_{4}$ $=-2c_{\beta}\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\beta_{k}\mathrm{Re}\langle\partial_{y}(iky\omega_{k}),NL_{k}\rangle\mathrm{d}k$
$=-2c_{\beta}\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\beta_{k}\mathrm{Re}\langle iky\partial_{y}\omega_{k},NL_{k}\rangle\mathrm{d}k-2c_{\beta}\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\beta_{k}\mathrm{Re}\langle ik\omega_{k},NL_{k}\rangle\mathrm{d}k$
@@ 3.5 Bound on $T_{4}$
For $T_{4}^{1}$, we note $|T_{4}^{1}|=|T_{3}|\lesssim\nu^{-\frac{2}{3}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}$. Since $\beta_{k}|k|\leq 1$, by the estimate of $T_{1}$, we have $|T_{4}^{2}|\lesssim\nu^{-\frac{2}{3}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}$.
@@ 3.6 Bound on $T_{5}$
and correspondingly $T_{5}\eqqcolon T_{5,LH,H}+T_{5,LH,L}+T_{5,HL,H^{\prime\prime}}+T_{5,HL,L^{\prime\prime}}$. For $T_{5,LH,H}$, we use Cauchy-Schwarz inequality in $y$, Young’s convolution inequality and Lemma 3.4 as before,
$|T_{5,LH,H}|$ $\lesssim\iint 1_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{J}\langle k\rangle^{m}\gamma_{k}\|y\omega_{k}\|_{2}\langle c\lambda_{k^{\prime}}t\rangle^{J}\langle k^{\prime}\rangle^{m}\|y\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}$
$\lesssim\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}^{2}\|y\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|y\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k$
$\lesssim\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}^{2}(\beta_{k}|k|^{2})^{-1}D_{k}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}(\gamma_{k}\nu)^{-1}D_{k}\mathrm{d}k\right)^{\frac{1}{2}}\mathcal{E}^{\frac{1}{2}}.$
@@
$|T_{5,LH,H}|\lesssim(\nu^{-1}\mathcal{D})^{\frac{1}{2}}(\nu^{-\frac{1}{3}}\mathcal{D})^{\frac{1}{2}}\mathcal{E}^{\frac{1}{2}}=\nu^{-\frac{2}{3}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
@@
$|T_{5,LH,L}|$ $\lesssim\iint 1_{|k-k^{\prime}|\lesssim\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{J}\langle k\rangle^{m}\nu^{-\frac{1}{3}}\gamma_{k}^{\frac{1}{2}}\|y\omega_{k}\|_{2}\langle c\lambda_{k^{\prime}}t\rangle^{J}\langle k^{\prime}\rangle^{m}\|y\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}$
@@
$\lesssim\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}\|y\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|y\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k$
$\lesssim\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}},$
where we have used (3.4) of Lemma 3.4. With the LH terms bounded, we now turn to estimate the HL terms, starting with $T_{5,HL,H^{\prime\prime}}$. We have $\gamma_{k}\lesssim\gamma_{k-k^{\prime}}$, since $|k|\lesssim|k-k^{\prime}|$. Thus
$|T_{5,HL,H^{\prime\prime}}|$ $\lesssim\iint 1_{|k-k^{\prime}|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{J}\langle k\rangle^{m}\gamma_{k}^{\frac{1}{2}}\|y\omega_{k}\|_{2}\langle c\lambda_{k-k^{\prime}}t\rangle^{J}\langle k-k^{\prime}\rangle^{m}\gamma_{k-k^{\prime}}^{\frac{1}{2}}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{\infty}\|y\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}\mathrm{d}k\mathrm{d}k^{\prime}$
$\lesssim\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}\|y\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}\|\nabla_{k}\psi_{k}\|_{\infty}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int\|y\nabla_{k}\omega_{k}\|_{2}\mathrm{d}k$
$\eqqcolon I_{1}\cdot I_{2}\cdot I_{3}.$
According to the definition of $\mathcal{E}$, we have $I_{1}\lesssim\mathcal{E}^{\frac{1}{2}}$. For $I_{2}$, we use Gagliardo-Nirenberg-Sobolev inequality, together with $\|k\nabla_{k}\psi_{k}\|_{2}\lesssim\|\Delta_{k}\psi_{k}\|_{2},\|\partial_{y}\nabla_{k}\psi_{k}\|\lesssim\|\Delta_{k}\psi_{k}\|_{2}$,
$I_{2}\lesssim\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}|k|^{-1}|k|\|\nabla_{k}\psi_{k}\|_{2}\|\partial_{y}\nabla_{k}\psi_{k}\|_{2}\mathrm{d}k\right)^{\frac{1}{2}}\lesssim\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}(\nu^{-\frac{1}{3}})^{-1}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\lesssim\nu^{-\frac{1}{3}}\mathcal{D}^{\frac{1}{2}}.$
By Hölder’s inequality and $\gamma_{k}\geq\nu^{-\frac{2}{3}}$, we obtain

### v4

@@ 3.1 Technical Lemmas
$\left(\int_{0}^{t}\left(\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\right)^{2}\mathrm{d}s\right)^{\frac{1}{2}}\lesssim\int_{\mathbb{R}}\left(\int_{0}^{t}\|\nabla_{k}\psi_{k}\|_{\infty}^{2}\mathrm{d}s\right)^{\frac{1}{2}}\mathrm{d}k\lesssim\nu^{-\frac{2}{3}}\mathcal{D}^{\frac{1}{2}}.$
Proof.
Using Gagliardo-Nirenberg-Sobolev inequality to estimate $\|\nabla_{k}\psi_{k}\|_{\infty}$,
$\|\nabla_{k}\psi_{k}\|_{\infty}\lesssim|k|^{-\frac{1}{2}}|k|^{\frac{1}{2}}\|\nabla_{k}\psi_{k}\|_{2}^{\frac{1}{2}}\|\partial_{y}\nabla_{k}\psi_{k}\|_{2}^{\frac{1}{2}}\lesssim|k|^{-\frac{1}{2}}\|\Delta_{k}\psi_{k}\|_{2}=|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}.$
Then applying Minkowski’s inequality for integrals, we have
$\left(\int_{0}^{t}\left(\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\right)^{2}\mathrm{d}s\right)^{\frac{1}{2}}\lesssim\int_{\mathbb{R}}\left(\int_{0}^{t}\|\nabla_{k}\psi_{k}\|_{\infty}^{2}\mathrm{d}s\right)^{\frac{1}{2}}\mathrm{d}k\lesssim\int_{\mathbb{R}}\left(\int_{0}^{t}|k|^{-1}\|\omega_{k}\|_{2}^{2}\mathrm{d}s\right)^{\frac{1}{2}}\mathrm{d}k.$
We split the integral domain into $|k|\geq\nu^{-\frac{1}{3}}$ and $|k|<\nu^{-\frac{1}{3}}$, then the $|k|\geq\nu^{-\frac{1}{3}}$ term obeys
$\int_{|k|\geq\nu^{-\frac{1}{3}}}\left(\int_{0}^{t}|k|^{-1}\|\omega_{k}\|_{2}^{2}\mathrm{d}s\right)^{\frac{1}{2}}\mathrm{d}k$ $=\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle k\rangle^{-m}(\gamma_{k}\nu)^{-\frac{1}{2}}|k|^{-\frac{1}{2}}\left(\int_{0}^{t}\langle k\rangle^{2m}\gamma_{k}\nu\|\omega_{k}\|_{2}^{2}\mathrm{d}s\right)^{\frac{1}{2}}\mathrm{d}k$
$\lesssim\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle k\rangle^{-2m}(\gamma_{k}\nu)^{-1}|k|^{-1}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\int_{0}^{t}\langle k\rangle^{2m}D_{k}\mathrm{d}s\mathrm{d}k\right)^{\frac{1}{2}}$
$\lesssim\nu^{-\frac{1}{6}}\mathcal{D}^{\frac{1}{2}},$
where we have used Hölder’s inequality. Then we integrate over $|k|<\nu^{-\frac{1}{3}}$.
$\int_{|k|<\nu^{-\frac{1}{3}}}\left(\int_{0}^{t}\left(|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}\right)^{2}\mathrm{d}s\right)^{\frac{1}{2}}\mathrm{d}k\lesssim\int_{|k|<\nu^{-\frac{1}{3}}}\nu^{-\frac{1}{2}}|k|^{-\frac{1}{2}}\mathrm{d}k\cdot\sup_{k\in\mathbb{R}}\left(\int_{0}^{t}\nu\|\omega_{k}\|_{2}^{2}\mathrm{d}s\right)^{\frac{1}{2}}\lesssim\nu^{-\frac{2}{3}}\mathcal{D}^{\frac{1}{2}}.$
Combine the two parts and then the lemma follows. ∎
Lemma 3.6.
The following estimate holds.
$\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}|k|\|\nabla_{k}\psi_{k}\|_{\infty}^{2}\mathrm{d}k\lesssim\nu^{-\frac{1}{3}}\widetilde{\mathcal{D}}.$
@@
$\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}|k|\|\nabla_{k}\psi_{k}\|_{\infty}^{2}\mathrm{d}k\lesssim\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}|k|\|\nabla_{k}\psi_{k}\|_{2}\|\partial_{y}\nabla_{k}\psi_{k}\|_{2}\mathrm{d}k\lesssim\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\|\omega_{k}\|_{2}^{2}\mathrm{d}k.$
@@ 3.2 Bound on $T_{1}$
$|T_{1,LH}|\lesssim\int_{0}^{t}\iint_{\mathbb{R}^{2}}\langle c\lambda_{k}s\rangle^{J}\langle k\rangle^{m}\|\omega_{k}\|_{2}\langle c\lambda_{k^{\prime}}s\rangle^{J}\langle k^{\prime}\rangle^{m}\|\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s,$
where we have used the property that $\lambda_{k}$ is monotonically increasing with respect to $|k|$. Then we use Young’s convolution inequality to place the $k$ and $k^{\prime}$ factors into $L^{2}$ and the $k-k^{\prime}$ factor into $L^{1}$,
$|T_{1,LH}|$ $\lesssim\int_{0}^{t}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\|\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}s$
$\lesssim\int_{0}^{t}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}(\gamma_{k}\nu)^{-1}D_{k}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\nu^{-1}D_{k}\mathrm{d}k\right)^{\frac{1}{2}}\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}s.$
@@
$|T_{1,LH}|\lesssim\int_{0}^{t}(\nu^{-\frac{1}{3}}\widetilde{\mathcal{D}})^{\frac{1}{2}}(\nu^{-1}\widetilde{\mathcal{D}})^{\frac{1}{2}}\mathcal{E}^{\frac{1}{2}}\mathrm{d}s\lesssim\nu^{-\frac{2}{3}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
Turning to $T_{1,HL}$, we similarly use Cauchy-Schwarz inequality in $y$, Young’s convolution inequality and Lemma 3.4,
$|T_{1,HL}|$ $\lesssim\int_{0}^{t}\iint_{\mathbb{R}^{2}}\langle c\lambda_{k}s\rangle^{J}\langle k\rangle^{m}\|\omega_{k}\|_{2}\langle c\lambda_{k-k^{\prime}}s\rangle^{J}\langle k-k^{\prime}\rangle^{m}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{2}\|\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
$\lesssim\int_{0}^{t}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\|\nabla_{k}\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{\mathbb{R}}\|\nabla_{k}\omega_{k}\|_{\infty}\mathrm{d}k\mathrm{d}s$
$\lesssim\int_{0}^{t}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}(\gamma_{k}\nu)^{-1}D_{k}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}^{-1}E_{k}\mathrm{d}k\right)^{\frac{1}{2}}\nu^{-\frac{5}{6}}\widetilde{\mathcal{D}}^{\frac{1}{2}}\mathrm{d}s$
$\lesssim\nu^{-\frac{2}{3}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
3.3 Bound on $T_{2},T_{3}$ and $T_{4}$
To estimate $T_{2}$, we apply integration by parts,
$T_{2}=-c_{\alpha}\int_{0}^{t}\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\alpha_{k}\mathrm{Re}\langle\Delta_{k}\omega_{k},NL_{k}\rangle\mathrm{d}k\mathrm{d}s.$
Using the same frequency decomposition as $T_{1}$, we note that $\alpha_{k}\leq\nu^{\frac{2}{3}}$, and the estimates are similar to $T_{1}$. The bound of $T_{2}$ is
$|T_{2}|\lesssim\nu^{-\frac{2}{3}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
For $T_{3}$, we still use the frequency decomposition $1=1_{|k-k^{\prime}|\leq\frac{|k|}{2}}+1_{|k-k^{\prime}|>\frac{|k|}{2}}$. Recalling the definition of $\beta_{k}$, we obtain $\beta_{k}|k|\leq 1$, and then, similarly to $T_{1}$, we have
$|T_{3}|\lesssim\nu^{-\frac{2}{3}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
We turn our attention now to $T_{4}$. We use integration by parts and separate $T_{4}$ into two parts,
$T_{4}$ $=-2c_{\beta}\int_{0}^{t}\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\beta_{k}\mathrm{Re}\langle\partial_{y}(iky\omega_{k}),NL_{k}\rangle\mathrm{d}k\mathrm{d}s$
$=-2c_{\beta}\int_{0}^{t}\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\beta_{k}\mathrm{Re}\langle iky\partial_{y}\omega_{k},NL_{k}\rangle\mathrm{d}k\mathrm{d}s-2c_{\beta}\int_{0}^{t}\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\beta_{k}\mathrm{Re}\langle ik\omega_{k},NL_{k}\rangle\mathrm{d}k\mathrm{d}s$
@@ 3.3 Bound on $T_{2},T_{3}$ and $T_{4}$
For $T_{4}^{1}$, we note
$|T_{4}^{1}|=|T_{3}|\lesssim\nu^{-\frac{2}{3}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
Since $\beta_{k}|k|\leq 1$, by the estimate of $T_{1}$, we have
$|T_{4}^{2}|\lesssim\nu^{-\frac{2}{3}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
@@ 3.4 Bound on $T_{5}$
and correspondingly $T_{5}\eqqcolon T_{5,LH,H}+T_{5,LH,L}+T_{5,HL,H^{\prime\prime}}+T_{5,HL,L^{\prime\prime}}$. For $T_{5,LH,H}$, we use $|k|\approx|k^{\prime}|$, Cauchy-Schwarz inequality in $y$, Young’s convolution inequality and Lemma 3.4 as before,
$|T_{5,LH,H}|$ $\lesssim\int_{0}^{t}\iint_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}s\rangle^{J}\langle k\rangle^{m}\gamma_{k}\|y\omega_{k}\|_{2}\langle c\lambda_{k^{\prime}}s\rangle^{J}\langle k^{\prime}\rangle^{m}\|y\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
$\lesssim\int_{0}^{t}\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}^{2}\|y\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\|y\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}s$
$\lesssim\int_{0}^{t}\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}^{2}(\beta_{k}|k|^{2})^{-1}D_{k}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}(\gamma_{k}\nu)^{-1}D_{k}\mathrm{d}k\right)^{\frac{1}{2}}\mathcal{E}^{\frac{1}{2}}\mathrm{d}s.$
@@
$|T_{5,LH,H}|\lesssim\int_{0}^{t}(\nu^{-1}\widetilde{\mathcal{D}})^{\frac{1}{2}}(\nu^{-\frac{1}{3}}\widetilde{\mathcal{D}})^{\frac{1}{2}}\mathcal{E}^{\frac{1}{2}}\mathrm{d}s\lesssim\nu^{-\frac{2}{3}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
@@
$|T_{5,LH,L}|$ $\lesssim\int_{0}^{t}\iint_{\mathbb{R}^{2}}\langle c\lambda_{k}s\rangle^{J}\langle k\rangle^{m}\nu^{-\frac{1}{3}}\gamma_{k}^{\frac{1}{2}}\|y\omega_{k}\|_{2}\langle c\lambda_{k^{\prime}}s\rangle^{J}\langle k^{\prime}\rangle^{m}\|y\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
@@
$\lesssim\int_{0}^{t}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}\|y\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\|y\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}s$
$\lesssim\nu^{-\frac{1}{2}}\int_{0}^{t}\mathcal{E}^{\frac{1}{2}}\widetilde{\mathcal{D}}^{\frac{1}{2}}\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}s.$
Then we use Hölder’s inequality in $s$ and Lemma 3.5,